# Zihao Zhang, Guodong Li, and Sihuang Hu

### v1

@@ Abstract
In this paper, we address this challenge for the case of two parity nodes (nk=2n-k=2) and sub-packetization $\ell=2$. We derive tight lower bounds on both the minimum repair bandwidth and the minimum I/O overhead. Furthermore, we present two explicit MDS array code constructions that achieve these bounds, respectively, offering practical code designs with provable repair efficiency.
@@ 1 Introduction
Fig. 1 illustrates an $(6,4)$ RS-coded stripe. In this example, if any two packets fail, the repair process requires transferring all four remaining packets, resulting in a repair bandwidth of $\mathcal{M}$ bits.
@@
Fig. 2 illustrates an $(6,4,8)$ MSR-coded stripe, where each packet is divided into $\ell=8$ sub-packets. In this example, if any single packet is lost, the repair process requires transferring four sub-packets from each of the remaining five packets; the resulting repair bandwidth is twenty sub-packets. Since each packet is divided into eight sub-packets, the repair bandwidth is equivalent to 2.5 packets, which significantly saves $37.5\%$ repair bandwidth compared to the (6, 4) RS-coded stripe in Fig. 1.
@@
The piggybacking framework [19] pioneered the construction of MDS codes with small sub-packetization and low repair bandwidth. This framework inherently retains the MDS property of RS codes while supporting arbitrary sub-packetization. Hitchhiker codes [20], a specialized piggybacking design with $\ell=2$, have been implemented in HDFS, reducing data packet repair bandwidth by 25%–45% compared to RS codes, though parity packet repair bandwidth remains unchanged. HashTag codes [21] similarly reduce repair bandwidth for data blocks. HashTag+ codes [22] further optimize repair bandwidth for both data and parity blocks. The Elastic Transformation [23] also converts RS codes into repair-efficient codes and supports configurable sub-packetization. Additionally, $\varepsilon$-MSR codes [24, 25] demonstrate that for sufficiently large $n$, logarithmic sub-packetization suffices to achieve $(1+\varepsilon)$-times the optimal repair bandwidth.
@@
Open Problem 1.
[15, Open Problem 9] Characterize the tradeoff between repair bandwidth, sub-packetization level, and field size for the general class of vector MDS codes.
@@ 1.1 Our contributions
We investigate the special case of Open Problem 1 under fixed parameters r=nk=2r=n-k=2 and $\ell=2$. Wu et al. addressed a variant of Open Problem 1 also in the special case of nk=2n-k=2 and $\ell=2$; their work requires the codes to be degraded-repair-friendly [32]. We impose no degraded-repair-friendly restriction. The redundancy $r=2$ is widely deployed in practice (e.g., RAID-6 [33] and Tencent Ultra-Cold Storage [34].) The sub-packetization $\ell=2$ further ensures that, within one stripe, every packet is accessed contiguously during repair, eliminating non-contiguous I/O.
@@
$\bar{\beta}(\mathcal{C})\geq\left\lceil\frac{5k}{4}\right\rceil.$
@@ 2 MDS codes and their repair schemes
where each $\mathbf{H}_{i}$ ($i\in[n]$) is an $r\ell\times\ell$ parity check sub-matrix over $\mathbb{F}_{q}$. Then, the $r$ nodes $C_{i_{1}},C_{i_{2}},\dots,C_{i_{r}}$ can be recovered from the remaining $k$ nodes if and only if the square matrix $\mathbf{H}_{i_{1}}\ \ \mathbf{H}_{i_{2}}\ \ \dots\ \ \mathbf{H}_{i_{r}}]$ is invertible.
@@
Repair scheme driven by a repair matrix. For any node $C_{i}$ of an $(n,k,\ell)$ MDS code $\mathcal{C}$, any linear repair process of $C_{i}$ can be described by an $\ell\times r\ell$ repair matrix $\mathbf{M}$ over $\mathbb{F}_{q}$. Conversely, any $\ell\times r\ell$ matrix $\mathbf{M}$ satisfying specific conditions can characterize a linear repair process for an MDS code $\mathcal{C}$.
@@
Let $\mathbf{M}$ be an $\ell\times r\ell$ matrix over $\mathbb{F}_{q}$. Multiplying $\mathbf{M}$ with the parity check equations (1), we obtain the following repair equations:
$\mathbf{M}\mathbf{H}_{1}C_{1}+\mathbf{M}\mathbf{H}_{2}C_{2}+\cdots+\mathbf{M}\mathbf{H}_{n}C_{n}=\bm{0}.$ (2)
@@
If the square matrix $\mathbf{M}\mathbf{H}_{i}$ is invertible for some $i$ (where $1\leq i\leq n$), then the repair equations (2) can be rewritten as
$C_{i}=-(\mathbf{M}\mathbf{H}_{i})^{-1}\sum_{j\in[n]\setminus\{i\}}\mathbf{M}\mathbf{H}_{j}C_{j},$
which means $C_{i}$ can be computed from the n1n-1 vectors $\mathbf{M}\mathbf{H}_{j}C_{j},j\in[n]\setminus\{i\}$.
@@
Repair bandwidth and repair I/O. Now we analyze the repair bandwidth and repair I/O of the repair process above, driven by the repair matrix $\mathbf{M}$. Suppose that the $\ell\times\ell$ matrix $\mathbf{M}\mathbf{H}_{i}$ is invertible for some $i\in[n]$. Then, each helper node $C_{j}$ ($j\in[n]\setminus\{i\}$) needs to transmit the vector $\mathbf{M}\mathbf{H}_{j}C_{j}$ to the failed node $C_{i}$. For each $j\in[n]$, we denote the rank of the $\ell\times\ell$ matrix $\mathbf{M}\mathbf{H}_{j}$ as $\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)$, and the number of nonzero columns in $\mathbf{M}\mathbf{H}_{j}$ as $\textnormal{nz}\left(\mathbf{M}\mathbf{H}_{j}\right)$. Then, the square matrix $\mathbf{M}\mathbf{H}_{j}$ can be decomposed as
$\mathbf{M}\mathbf{H}_{j}=\mathbf{L}_{j}\cdot\mathbf{R}_{j},$
@@
where $\mathbf{L}_{j}$ is an $\ell\times\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)$ matrix and $\mathbf{R}_{j}$ is an $\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)\times\ell$ matrix. We can obtain $\mathbf{L}_{j}$ by selecting $\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)$ independent columns of $\mathbf{M}\mathbf{H}_{j}$, and obtain $\mathbf{R}_{j}$ from the linear combination coefficients for the columns of $\mathbf{L}_{j}$ to generate the columns in $\mathbf{M}\mathbf{H}_{j}$. We have $\textnormal{rank}\left(\mathbf{R}_{j}\right)=\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)$, $\textnormal{nz}\left(\mathbf{R}_{j}\right)=\textnormal{nz}\left(\mathbf{M}\mathbf{H}_{j}\right)$, and the repair equations (2) can be further rewritten as
$C_{i}=-(\mathbf{M}\mathbf{H}_{i})^{-1}\cdot\sum_{j\in[n]\setminus\{i\}}\mathbf{L}_{j}(\mathbf{R}_{j}C_{j}).$
@@
Since $\mathbf{L}_{j}$, $j\in[n]\setminus\{i\}$, can be calculated from $\mathbf{M}$ and the parity check sub-matrix $\mathbf{H}_{j}$, it is sufficient to transmit the n1n-1 vectors $\mathbf{R}_{j}C_{j}$, $j\in[n]\setminus\{i\}$ to the failed node $C_{i}$ for the repair.
@@
Note that each $\mathbf{R}_{j}C_{j}$ is a vector of length $\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)$ over $\mathbb{F}_{q}$, and transmitting $\mathbf{R}_{j}C_{j}$ requires to read $\textnormal{nz}\left(\mathbf{R}_{j}\right)$ symbols from $C_{j}$. Therefore, during the repair process of $C_{i}$ driven by the repair matrix $\mathbf{M}$, the repair bandwidth, i.e., the total number of symbols transmitted from the helper nodes to the failed node, is
$\operatorname{BW}(\mathbf{M})=\sum_{j\in[n]\setminus\{i\}}\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)=\sum_{j\in[n]}\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)-\ell.$ (3)
@@
$\operatorname{IO}(\mathbf{M})$ $=\sum_{j\in[n]\setminus\{i\}}\textnormal{nz}\left(\mathbf{R}_{j}\right)=\sum_{j\in[n]\setminus\{i\}}\textnormal{nz}\left(\mathbf{M}\mathbf{H}_{j}\right)$ (4)
$=\sum_{j\in[n]}\textnormal{nz}\left(\mathbf{M}\mathbf{H}_{j}\right)-\ell.$
@@
$d(\mathbf{M})=\left|\{i\in[n]:\mathbf{M}\mathbf{H}_{i}\neq\mathbf{O}\}\right|-1.$
@@
For an $(n,k,\ell)$ MDS code $\mathcal{C}$ and a nonzero matrix $\mathbf{M}$ of size $\ell\times r\ell$, we have $d(\mathbf{M})\geq k$.
@@
If $d(\mathbf{M})\leq k-1$, then there exist $r$ distinct $a_{1},\dots,a_{r}\in[n]$ such that $\mathbf{M}\cdot[\mathbf{H}_{a_{1}},\mathbf{H}_{a_{2}},\cdots,\mathbf{H}_{a_{r}}]=\mathbf{O}$. By the MDS property, $\mathbf{H}_{a_{1}},\mathbf{H}_{a_{2}},\cdots,\mathbf{H}_{a_{r}}]$ is invertible, and hence $\mathbf{M}$ is a zero matrix, which leads to a contradiction. ∎
We denote the index set of nodes which can be repaired by an $\ell\times r\ell$ matrix $\mathbf{M}$ as
$\mathcal{R}(\mathbf{M}):=\{i\in[n]:\ \textnormal{rank}\left(\mathbf{M}\mathbf{H}_{i}\right)=\ell\}.$ (5)
@@
$\mathcal{M}_{i}:=\left\{\mathbf{M}\in\mathbb{F}_{q}^{\ell\times r\ell}:i\in\mathcal{R}(\mathbf{M})\right\}.$ (6)
@@
$\bar{\beta}(\mathcal{C})$ $=\frac{1}{n}\sum_{i\in[n]}\beta_{i}=\frac{1}{n}\sum_{i\in[n]}\min_{\mathbf{M}\in\mathcal{M}_{i}}\{\operatorname{BW}(\mathbf{M})\},$ (7)
$\beta(\mathcal{C})$ $=\max_{i\in[n]}\{\beta_{i}\}=\max_{i\in[n]}\min_{\mathbf{M}\in\mathcal{M}_{i}}\{\operatorname{BW}(\mathbf{M})\},$ (8)
$\bar{\gamma}(\mathcal{C})$ $=\frac{1}{n}\sum_{i\in[n]}\gamma_{i}=\frac{1}{n}\sum_{i\in[n]}\min_{\mathbf{M}\in\mathcal{M}_{i}}\{\operatorname{IO}(\mathbf{M})\},$ (9)
$\gamma(\mathcal{C})$ $=\max_{i\in[n]}\{\beta_{i}\}=\max_{i\in[n]}\min_{\mathbf{M}\in\mathcal{M}_{i}}\{\operatorname{IO}(\mathbf{M})\},.$ (10)
@@ 3 Fundamental Limits
This section establishes the lower bounds of 1) the avg-min repair bandwidth (7), and 2) the avg-min repair I/O (9) of a $(k+2,k,2)$ MDS array code $\mathcal{C}$ over finite field $\mathbb{F}_{q}$, where $r=n-k=2,\ell=2$ and the parity check sub-matrix $\mathbf{H}_{i}$ for node $C_{i}$ (where $i\in[n]$) is of size $4\times 2$. By Lemma 5, we have $d(\mathbf{M})\in\{k,k+1\}$ for any nonzero repair matrix $\mathbf{M}$ of size $2\times 4$. The following lemma shows that we only need to consider the case of $d(\mathbf{M})=k+1$ when we measure the lower bounds on avg-min repair bandwidth and avg-min repair I/O.
@@
For each node $C_{i}$ of an $(k+2,k,2)$ MDS array code, there exists $\mathbf{M}\in\mathcal{M}_{i}$ of $d(\mathbf{M})=k+1$ such that $\operatorname{IO}(\mathbf{M})\leq 2k$, and therefore $\operatorname{BW}(\mathbf{M})\leq 2k$.
@@
Without loss of generality, we assume $i=1$ and parity check matrix
$\mathbf{H}=\begin{bmatrix}\mathbf{A}_{1}&\dots&\mathbf{A}_{k}&\mathbf{I}_{2}&\mathbf{O}\\ \mathbf{B}_{1}&\dots&\mathbf{B}_{k}&\mathbf{O}&\mathbf{I}_{2}\end{bmatrix}$.
Let $\mathbf{M}_{1}=\begin{bmatrix}1&0&&\\ &&1&0\end{bmatrix}$ and $\mathbf{M}_{2}=\begin{bmatrix}1&0&&\\ &&0&1\end{bmatrix}$. As $\mathbf{A}_{i},\mathbf{B}_{i},i\in[k]$ are all invertible matrices, it’s obvious that $d(\mathbf{M}_{1})=d(\mathbf{M}_{2})=k+1$ and $\operatorname{IO}(\mathbf{M}_{1}),\operatorname{IO}(\mathbf{M}_{2})\leq 2k$. Now we claim that $\mathbf{M}_{1}\in\mathcal{M}_{1}$ or $\mathbf{M}_{2}\in\mathcal{M}_{1}$.
If $\mathbf{M}_{1}\notin\mathcal{M}_{1}$ and $\mathbf{M}_{2}\notin\mathcal{M}_{1}$, then
$\textnormal{rank}\left(\mathbf{M}_{1}\begin{bmatrix}\mathbf{A}_{1}\\ \mathbf{B}_{1}\end{bmatrix}\right)=\textnormal{rank}\left(\begin{bmatrix}\mathbf{e}_{1}\mathbf{A}_{1}\\ \mathbf{e}_{1}\mathbf{B}_{1}\end{bmatrix}\right)=1$
and
$\textnormal{rank}\left(\mathbf{M}_{2}\begin{bmatrix}\mathbf{A}_{1}\\ \mathbf{B}_{1}\end{bmatrix}\right)=\textnormal{rank}\left(\begin{bmatrix}\mathbf{e}_{1}\mathbf{A}_{1}\\ \mathbf{e}_{2}\mathbf{B}_{1}\end{bmatrix}\right)=1$
where $\mathbf{e}_{1}=\begin{pmatrix}1&0\end{pmatrix}$ and $\mathbf{e}_{2}=\begin{pmatrix}0&1\end{pmatrix}$. Therefore, we have $\langle\mathbf{e}_{1}\mathbf{B}_{1}\rangle=\langle\mathbf{e}_{1}\mathbf{A}_{1}\rangle=\langle\mathbf{e}_{2}\mathbf{B}_{1}\rangle$. This contradicts the fact that $\mathbf{B}_{1}$ is invertible.
In conclusion, there exists $\mathbf{M}\in\mathcal{M}_{i}$ of $d(\mathbf{M})=k+1$ such that $\operatorname{IO}(\mathbf{M})\leq 2k$. ∎
Definition 1 describes a total order on $2^{[n]}$. It first compares subsets by their cardinality. Then, for subsets of the same cardinality, it compares them by the dictionary order.

### v2

@@ Abstract
In this paper, we address this challenge for the case of two parity nodes (nk=2n-k=2) and sub-packetization $\ell=2$. Under these parameters, we establish a correspondence between repair schemes and point sets on the projective line $\mathbb{P}^{1}$, and then derive a lower bound on repair bandwidth utilizing the sharply 3-transitive action of $\text{PGL}_{2}(\mathbb{F}_{q})$. Furthermore, we extend this lower bound to the repair I/O, and construct two classes of explicit MDS array codes that achieve these bounds, offering practical code designs with provable repair efficiency.
@@ 1 Introduction
Fig. 1 illustrates a $(6,4)$ RS-coded stripe. In this example, if any single packet fails, the repair process requires transferring all four remaining packets, resulting in a repair bandwidth of $\mathcal{M}$ bits.
@@
Fig. 2 illustrates a $(6,4,8)$ MSR-coded stripe, where each packet is divided into $\ell=8$ sub-packets. In this example, if any single packet is lost, the repair process requires transferring four sub-packets from each of the remaining five packets; the resulting repair bandwidth is twenty sub-packets. Since each packet is divided into eight sub-packets, the repair bandwidth is equivalent to 2.5 packets, which significantly saves $37.5\%$ repair bandwidth compared to the (6, 4) RS-coded stripe in Fig. 1.
@@
The piggybacking framework [30] pioneered the construction of MDS codes with small sub-packetization and low repair bandwidth. This framework inherently retains the MDS property of RS codes while supporting arbitrary sub-packetization. Hitchhiker codes [28], a specialized piggybacking design with $\ell=2$, have been implemented in HDFS, reducing data packet repair bandwidth by 25%–45% compared to RS codes, though parity packet repair bandwidth remains unchanged. HashTag codes [14] similarly reduce repair bandwidth for data blocks. HashTag+ codes [15] further optimize repair bandwidth for both data and parity blocks. The Elastic Transformation [33] also converts RS codes into repair-efficient codes and supports configurable sub-packetization. Additionally, $\varepsilon$-MSR codes [31, 26] demonstrate that for sufficiently large $n$, logarithmic sub-packetization suffices to achieve $(1+\varepsilon)$-times the optimal repair bandwidth. Other notable constructions with near-optimal bandwidth are proposed in [18, 17, 5, 13].
@@
Open Problem.
[25, Open Problem 9] Characterize the tradeoff between repair bandwidth, sub-packetization level, and field size for the general class of vector MDS codes.
@@ 1.1 Our Contributions
We investigate the special case of Open Problem Open Problem under fixed parameters r=nk=2r=n-k=2 and $\ell=2$. Wu et al. addressed a variant of Open Problem Open Problem also in the special case of nk=2n-k=2 and $\ell=2$; their work requires the codes to be degraded-repair-friendly [37]. We impose no degraded-repair-friendly restriction. The redundancy $r=2$ is widely deployed in practice (e.g., RAID-6 [4] and Tencent Ultra-Cold Storage [12].) The sub-packetization $\ell=2$ further ensures that, within one stripe, every packet is accessed contiguously during repair, eliminating non-contiguous I/O.
@@
$\beta(\mathcal{C})\geq\left\lceil\frac{5k}{4}\right\rceil.$
@@ 2 MDS Codes and Their Repair Schemes
where each $H_{i}$ ($i\in[n]$) is an $r\ell\times\ell$ parity check sub-matrix over $\mathbb{F}_{q}$. Then, the $r$ nodes $C_{i_{1}},C_{i_{2}},\dots,C_{i_{r}}$ can be recovered from the remaining $k$ nodes if and only if the square matrix $H_{i_{1}}\ \ H_{i_{2}}\ \ \dots\ \ H_{i_{r}}]$ is invertible.
@@
Repair scheme driven by a repair matrix. For any node $C_{i}$ of an $(n,k,\ell)$ MDS code $\mathcal{C}$, any linear repair process of $C_{i}$ can be described by an $\ell\times r\ell$ repair matrix $M$ over $\mathbb{F}_{q}$. Conversely, any $\ell\times r\ell$ matrix $M$ satisfying specific conditions can characterize a linear repair process for an MDS code $\mathcal{C}$.
@@
Let $M$ be an $\ell\times r\ell$ matrix over $\mathbb{F}_{q}$. Multiplying $M$ with the parity check equations (1), we obtain the following repair equations:
$MH_{1}C_{1}+MH_{2}C_{2}+\cdots+MH_{n}C_{n}=\bm{0}.$ (2)
@@
If the square matrix $MH_{i}$ is invertible for some $i$ (where $1\leq i\leq n$), then the repair equations (2) can be rewritten as
$C_{i}=-(MH_{i})^{-1}\sum_{j\in[n]\setminus\{i\}}MH_{j}C_{j},$
which means $C_{i}$ can be computed from the n1n-1 vectors $MH_{j}C_{j},j\in[n]\setminus\{i\}$.
@@
Repair bandwidth and repair I/O. Now we analyze the repair bandwidth and repair I/O of the repair process above, driven by the repair matrix $M$. Suppose that the $\ell\times\ell$ matrix $MH_{i}$ is invertible for some $i\in[n]$. Then, each helper node $C_{j}$ ($j\in[n]\setminus\{i\}$) needs to transmit the vector $MH_{j}C_{j}$ to the failed node $C_{i}$. For each $j\in[n]$, we denote the rank of the $\ell\times\ell$ matrix $MH_{j}$ as $\textnormal{rank}\left(MH_{j}\right)$, and the number of nonzero columns in $MH_{j}$ as $\textnormal{nz}\left(MH_{j}\right)$. Then, the square matrix $MH_{j}$ can be decomposed as
$MH_{j}=L_{j}\cdot R_{j},$
@@
where $L_{j}$ is an $\ell\times\textnormal{rank}\left(MH_{j}\right)$ matrix and $R_{j}$ is an $\textnormal{rank}\left(MH_{j}\right)\times\ell$ matrix. We can obtain $L_{j}$ by selecting $\textnormal{rank}\left(MH_{j}\right)$ independent columns of $MH_{j}$, and obtain $R_{j}$ from the linear combination coefficients for the columns of $L_{j}$ to generate the columns in $MH_{j}$. We have $\textnormal{rank}\left(R_{j}\right)=\textnormal{rank}\left(MH_{j}\right)$, $\textnormal{nz}\left(R_{j}\right)=\textnormal{nz}\left(MH_{j}\right)$, and the repair equations (2) can be further rewritten as
$C_{i}=-(MH_{i})^{-1}\cdot\sum_{j\in[n]\setminus\{i\}}L_{j}(R_{j}C_{j}).$
@@
Since $L_{j}$, $j\in[n]\setminus\{i\}$, can be calculated from $M$ and the parity check sub-matrix $H_{j}$, it is sufficient to transmit the n1n-1 vectors $R_{j}C_{j}$, $j\in[n]\setminus\{i\}$ to the failed node $C_{i}$ for the repair.
@@
Note that each $R_{j}C_{j}$ is a vector of length $\textnormal{rank}\left(MH_{j}\right)$ over $\mathbb{F}_{q}$, and transmitting $R_{j}C_{j}$ requires to read $\textnormal{nz}\left(R_{j}\right)$ symbols from $C_{j}$. Therefore, during the repair process of $C_{i}$ driven by the repair matrix $M$, the repair bandwidth, i.e., the total number of symbols transmitted from the helper nodes to the failed node, is
$\operatorname{BW}(M)=\sum_{j\in[n]\setminus\{i\}}\textnormal{rank}\left(MH_{j}\right)=\sum_{j\in[n]}\textnormal{rank}\left(MH_{j}\right)-\ell.$ (3)
@@
$\operatorname{IO}(M)$ $=\sum_{j\in[n]\setminus\{i\}}\textnormal{nz}\left(R_{j}\right)=\sum_{j\in[n]\setminus\{i\}}\textnormal{nz}\left(MH_{j}\right)$ (4)
$=\sum_{j\in[n]}\textnormal{nz}\left(MH_{j}\right)-\ell.$
@@
$d(M)=\left|\{i\in[n]:MH_{i}\neq O\}\right|-1.$
@@
For an $(n,k,\ell)$ MDS code $\mathcal{C}$ and a nonzero matrix $M$ of size $\ell\times r\ell$, we have $d(M)\geq k$.
@@
If $d(M)\leq k-1$, then there exist $r$ distinct $a_{1},\dots,a_{r}\in[n]$ such that $M\cdot[H_{a_{1}},H_{a_{2}},\cdots,H_{a_{r}}]=\mathbf{O}$. By the MDS property, $H_{a_{1}},H_{a_{2}},\cdots,H_{a_{r}}]$ is invertible, and hence $M$ is a zero matrix, which leads to a contradiction. ∎
We denote the index set of nodes which can be repaired by an $\ell\times r\ell$ matrix $M$ as
$\mathcal{R}(M):=\{i\in[n]:\ \textnormal{rank}\left(MH_{i}\right)=\ell\}.$ (5)
@@
$\mathcal{M}_{i}:=\left\{M\in\mathbb{F}_{q}^{\ell\times r\ell}:i\in\mathcal{R}(M)\right\}.$ (6)
@@
$\bar{\beta}(\mathcal{C})$ $=\frac{1}{n}\sum_{i\in[n]}\beta_{i}=\frac{1}{n}\sum_{i\in[n]}\min_{M\in\mathcal{M}_{i}}\{\operatorname{BW}(M)\},$ (7)
$\beta(\mathcal{C})$ $=\max_{i\in[n]}\{\beta_{i}\}=\max_{i\in[n]}\min_{M\in\mathcal{M}_{i}}\{\operatorname{BW}(M)\},$ (8)
$\bar{\gamma}(\mathcal{C})$ $=\frac{1}{n}\sum_{i\in[n]}\gamma_{i}=\frac{1}{n}\sum_{i\in[n]}\min_{M\in\mathcal{M}_{i}}\{\operatorname{IO}(M)\},$ (9)
$\gamma(\mathcal{C})$ $=\max_{i\in[n]}\{\beta_{i}\}=\max_{i\in[n]}\min_{M\in\mathcal{M}_{i}}\{\operatorname{IO}(M)\},.$ (10)
@@ 3 Fundamental Limits
This section establishes the lower bounds of 1) the avg-min repair bandwidth (7), and 2) the avg-min repair I/O (9) of a $(k+2,k,2)$ MDS array code $\mathcal{C}$ over finite field $\mathbb{F}_{q}$, where $r=n-k=2,\ell=2$ and the parity check sub-matrix $H_{i}$ for node $C_{i}$ (where $i\in[n]$) is of size $4\times 2$. Throughout this section, we work with such codes with a fixed parity-check matrix in systematic form
$H=\begin{bmatrix}A_{1}&\cdots&A_{k}&I_{2}&O\\ B_{1}&\cdots&B_{k}&O&I_{2}\end{bmatrix}$,
where each block $A_{i},B_{i}\in\mathbb{F}_{q}^{2\times 2}$ is invertible by the MDS property.
@@
For a $(k+2,k,2)$ MDS array code, to determine the minimum repair bandwidth or repair I/O of any node $C_{i}$, it suffices to consider repair matrices with repair degree $d(M)=k+1$.
@@
For any $(n,k,\ell)$ MDS code, the repair degree $d(M)$ must satisfy $d(M)\geq k$. Since $n=k+2$, the repair degree is constrained to $d(M)\in\{k,k+1\}$. Note that trivial repair ($d(M)=k$) yields the fixed bandwidth and I/O of $2k$. Therefore, it is sufficient to demonstrate the existence of a repair matrix $M$ for $C_{i}$ such that $d(M)=k+1$ and both repair bandwidth and repair I/O $\leq 2k$.
Without loss of generality, assume $C_{1}$ fails. Consider the matrices:
$M_{1}=\begin{bmatrix}1&0&0&0\\ 0&0&1&0\end{bmatrix}\quad\text{and}\quad M_{2}=\begin{bmatrix}1&0&0&0\\ 0&0&0&1\end{bmatrix}$.
Since $W_{1}$ is invertible, at least one of $M_{1}H_{1}$ or $M_{2}H_{1}$ is invertible. Thus, there exists a matrix $M\in\{M_{1},M_{2}\}$ capable of repairing $C_{1}$ with $d(M)=k+1$ and repair I/O $\operatorname{IO}(M)\leq 2k$. Since $\operatorname{BW}(M)\leq\operatorname{IO}(M)$, its repair bandwidth also more than $2k$.
This confirms that the minimum bandwidth and I/O are determined within the case $d(M)=k+1$. ∎
Based on Lemma 6, we assume that the repair degree is $k+1$ in the subsequent analysis.
To analyze the minimum repair bandwidth and repair I/O, we define a total order on the power set $2^{[n]}$. It first compares subsets by their cardinality. Then, for subsets of the same cardinality, it compares them by the dictionary order.